OpenStudy (anonymous):

WILL FAN AND MEDAL! Word problems are my weakness.. 7. If an object is dropped from a height of 200 feet, the function h(t)=-16t^2+200 gives the height of the objects after t seconds. Approximately, when will the object hit the ground? A.200.00 seconds B.184.00 seconds C.3.54 seconds D.0.78 seconds

OpenStudy (anonymous):

8. A ball is thrown in the air with an upward velocity of 28ft/s. Its height (h) in feet after t seconds is given by the function h=-16t^2+28t+7. How long does it take the ball to reach its maximum height? What is the ball's maximum height? Round to the nearest hundreth, if necessary. A. Reaches a maximum height if 7 feet after 1.75 seconds B. Reaches a maximum height of 43.75 feet after 0.88 seconds C. Reaches a maximum height of 17.5 feet after 0.88 seconds D. Reaches a maximum height of 19.25 feet after 0.88 seconds

OpenStudy (alekos):

Make h = 0 and solve for t

OpenStudy (anonymous):

13. A catapult launches a boulder with an up ward velocity of 184 feet per second. The height of the boulder,(h), in feet after t seconds is given by the function h(t)=-16t^2+184t+20. How long does it take the boulder to reach its maximum height? Round to the nearest hundredth, if necessary. A. Reaches a maximum height of 11.6 feet after 5.75 seconds B. Reaches a maximum height of 549 feet after 11.5 seconds C. Reaches a maximum height of 549 feet after 5.75 seconds D. Reaches a maximum height of 23.2 feet after 11.6 seconds

OpenStudy (anonymous):

@alekos can u take me through it step by step because like i said word problems are my weakness and i don't understand

OpenStudy (alekos):

0 = -16t^2 + 200 Now solve for t

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@alekos is it C?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@jakashaka123 do you know? @iGreen do you know?

OpenStudy (anonymous):

im sorry i dont =(

OpenStudy (anonymous):

@Michele_Laino @dan815

OpenStudy (anonymous):

@jakashaka123 thats okay!

OpenStudy (michele_laino):

question 7) If I call with \tau the time when ball hits the ground, then we have to request that the subsequent condition holds: \[\Large h\left( \tau \right) = 0\] or: \[ \Large - 16{\tau ^2} + 200 = 0\] so what is \tau?

OpenStudy (anonymous):

@Michele_Laino i thought 7 is C

OpenStudy (anonymous):

okay so i did 7.C 8.D and 13 i need help with

OpenStudy (anonymous):

@iGreen @dan815

OpenStudy (anonymous):

#1=3.54sec because 16t^2=200

OpenStudy (anonymous):

yeah @jamaica25 i got that one but i need help with the third one

OpenStudy (anonymous):

#2 19.26 ft after 0.88sec because time =-b/2a =-28/2*16=0.88 sub in equation h=19.26

OpenStudy (anonymous):

@jamaica25 i got that one too i need the third one

OpenStudy (anonymous):

if you follow my answer the third is the same reasoning

OpenStudy (anonymous):

can u like walk me through the problem so i understand it better? @jamaica25

OpenStudy (anonymous):

use the formula -b/2a =-184/(2*16) =5.74 sec.

OpenStudy (anonymous):

so the third one is either a or c? @jamaica25

OpenStudy (anonymous):

yes can you use the formula?

OpenStudy (anonymous):

yes? @jamaica25

OpenStudy (anonymous):

put 5.74 into the height formula to get which height

OpenStudy (anonymous):

a? @jamaica25

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:( c? @jamaica25

OpenStudy (anonymous):

-16(5.75)^2+184(5.74)+20 =549

OpenStudy (anonymous):

ooh i see now

OpenStudy (anonymous):

the answers are not important understanding the formula is

OpenStudy (anonymous):

yeah i was really trying to understand the formula for this problem but these types of problems are my weakness in math

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you so much for all your help!

OpenStudy (anonymous):

@jamaica25 i have one more like this do you think you can help?